OpenStudy (anonymous):

Solve. x2 + x – 6 = 0 A. {2, 3} B. {2, –3} C. {3, –2} D. {6, –1}

OpenStudy (anonymous):

@Michele_Laino

pooja195 (pooja195):

Do we need to use the quadratic formula again or can we factor?

OpenStudy (anonymous):

um, it doesnt say!

pooja195 (pooja195):

ok lets just factor 

pooja195 (pooja195):

what multiplies to -6 but adds up to -1?

pooja195 (pooja195):

facotor out -6 if it helps :)

pooja195 (pooja195):



OpenStudy (anonymous):

oh okay

pooja195 (pooja195):

(x-3)=0 (x+2)=)

pooja195 (pooja195):

x-3=0 x+2=0 solve to get the answers :)

OpenStudy (anonymous):

3 and -2

pooja195 (pooja195):

yep

OpenStudy (anonymous):

Oh Jee use quadratic equation formula... x2 + x – 6 = 0 => ax^2 + bx + c =o \[\frac{-b \pm \sqrt{b ^{2}-4ac} }{ 2a}\]

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

x = -3, x = 2

pooja195 (pooja195):

yes

OpenStudy (anonymous):

okay! i will post another question!

pooja195 (pooja195):

:)